# Relative Distance Guided Dynamic Partition Learning for Scale-Invariant UAV-View Geo-Localization

### v2

@@ Abstract
UAV-view Geo-Localization (UVGL) presents substantial challenges, particularly due to the disparity in visual appearance between drone-captured imagery and satellite perspectives. Existing methods usually assume consistent scaling factor across different views. Therefore, they adopt predefined partition alignment and extract viewpoint-invariant representation by constructing a variety of part-level features. However, the scaling assumption is not always hold in the real-world scenarios that variations of UAV flight state leads to the scale mismatch of cross-views, resulting in serious performance degradation. To overcome this issue, we propose a partition learning framework based on relative distance, which alleviates the dependence on scale consistency while mining fine-grained features. Specifically, we propose a distance guided dynamic partition learning strategy (DGDPL), consisting of a square partition strategy and a distance-guided adjustment strategy. The former is utilized to extract fine-grained features and global features in a simple manner. The latter calculates the relative distance ratio between drone- and satellite-view to adjust the partition size, thereby explicitly aligning the semantic information between partition pairs. Furthermore, we propose a saliency-guided refinement strategy to refine part-level features, so as to further improve the retrieval accuracy. Extensive experiments show that our approach achieves superior geo-localization accuracy across various scale-inconsistent scenarios, and exhibits remarkable robustness against scale variations. The code will be released.
@@
image retrieval, geo-localization, dynamic partition, relative distance ratio, scale variations
@@ I Introduction
Unmanned Aerial Vehicles (UAVs) have gained popularity due to their ability to efficiently capture data with few occlusions and rich content. Differ from ground-views, UAVs can easily acquire multi-resolution images by changing the flight altitude and camera parameters. Benefiting from these advantages, UAVs play an irreplaceable role in various fields, including agricultural operations [1], automatic driving [2] and aerial photography [3]. Regardless of the applications, precise positioning and navigation of UAVs are indispensable, which mainly rely on a global navigation satellite system (GNSS) and high-quality communication environment [4, 5, 6]. In scenarios with weak or failed signals, GNSS will incorrectly determine the geographic localization of UAVs. Therefore, vision-based UAV-view geo-localization (UVGL) develops a meaningful hot research, aiming to release the dependence on GNSS. As the key of this task, cross-view matching algorithms calculates the similarity between drone- and satellite-view images to achieve image matching. However, due to the appearance difference such as scale and color between drone-views and satellite-views, traditional image matching algorithms [7, 8, 9] are imprecise.
@@
Figure 1: The simplified diagram of our research motivation. Given drone-view images captured from different altitudes (high-yellow box, middle-green box, low-blue box) and the matched satellite-view image, part-level image representations are constructed using a fixed partition strategy and our dynamic partition strategy, respectively. As can be observed, the extended/indented partition pairs have more consistent semantic content.
@@
With the rapid development of deep learning, current methods [10, 11, 12, 13, 14, 15, 16, 17, 18] utilize pre-trained backbone networks to extract high-level semantic representations and perform image matching based on feature distances, achieving superior accuracy than non-learning-based algorithms. To obtain robust discriminative representations, plentiful partition learning strategies have been proposed, which constructs fine-grained features by segmenting high-level features to fully exploit contextual information. Depending on the pattern of feature segmentation, partition learning can be divided into two categories: soft-partition learning [19, 20, 21, 22, 23] and hard-partition learning [24, 25, 26, 27, 28, 29, 30]. The former classifies features based on feature values to distinguish the categories of semantic information, so the shape of partitions is uncontrollable. FSRA [19], as a representative of soft-partition learning, confirms that fine-grained features have a consistent distribution with image content and can boost model performance. In contrast, inspired by the distribution similarity between drone- and satellite-views, hard-partition strategies set predefined templates with fixed shape to divide high-level features. As the pioneering hard-partition algorithm, LPN [24] designs non-overlapping square-ring templates to build fine-grained features. Subsequently, several studies [31, 32, 33] further optimize the shape and number of templates to facilitate discriminative representation. Benefiting from the explicit utilization of contextual information, hard-partition strategies reveal superior accuracy and compatibility than soft-partition strategies under same settings. Nevertheless, hard-partition strategies are more sensitive to spatial changes of input images. During the data collection process, the variations of UAV flight state will lead to the scale mismatch of cross-views. The predefined segmentation templates constrain misaligned semantic information to same parts, thus interfering with feature matching.
@@
As shown in Fig. 1, we depict the limitation of hard-partition strategies in dealing with scale-misaligned image pairs. Taking LPN [24] as an example, the partition pairs contain similar semantic information in the case of similar cross-view scales (see drone-view with green box). When UAV altitude changes, partition pairs of fixed shape contain misaligned semantic information, defined as imperfect matching (see drone-view with yellow or blue box). Following the imaging law of ”objects appearing larger when near and smaller when distant” in photography, we argue that adjusting the partition size based on the shooting distance of the UAV can ensure the content consistency of partition pairs, thus boosting the matching of cross-view partition pairs. Concretely, for high-altitude (i.e., long distance) UAV-view, indent templates can reduce redundancy in the part-level representations, while expanded templates contributes to the content supplement of partitions for low-altitude (i.e., short distance) UAV-view. Imitating the process mentioned above, we propose a distance guided dynamic partition learning framework named DGDPL, to enhance the content consistency of partition pairs when the shooting distance of UAV changed. Specifically, we design a square partition strategy (SPS) as shown in Case(II) of Fig. 2, which has two typical advantages: 1) multiple partitions constructed by SPS emphasize both fine-grained and global features, thereby exhibiting robustness against scale variations; 2) solid partitions are readily generated and have strong compatibility with subsequent algorithms. To promote the content consistency of partition pairs, we further propose a distance-guided adjustment strategy (DGAS). We construct a scale factor $\beta$ to represent the relative distance of drone- and satellite-view, and inject it into SPS to dynamically adjust the partition size of drone-views. As shown in Case(I) and Case(III) of Fig. 2, our DGAS controls variations of partitions, both in direction and degree. Furthermore, we design a saliency-guided refinement strategy (SGRS) to identify targets and environments based on feature activation degrees. The SGRS maintains robustness to scale changes while improving model matching accuracy. Extensive experiments on University-1652 show that our DGDPL framework exhibits excellent robustness to cross-view scale changes, and achieves remarkable performance in scenarios with inconsistent cross-view scales. For instance, the average R@1 of our DGDPL on multiple scale scenarios is nearly 30.70% and 16.64% higher than that of baseline and LPN, respectively.
Our main contributions are as follows:
  • We propose DGDPL, including a square partition strategy and a distance-guided adjustment strategy, to achieve accurate UAV-view geo-localization against scale variations of drone-views.
  • We formulate a saliency-guided refinement strategy to refine part-level features, thus improving retrieval accuracy while maintaining robustness to scale changes.
@@
  • Extensive experiments show that, in comparison to both hard-partition and soft-partition strategies, our method achieves splendid retrieval performance in scenes with inconsistent cross-view scales.
@@
Figure 2: Distance guided dynamic partition learning strategy ($N=3$ for illustration). Case(II): the default square partition process, i.e., the partition size of satellite- and drone-view is same. Case(I): the partitions are indented based on scale factor $\theta_{1}(>0)$ to fit the drone-view with long distance. Case(III): the partitions are expended based on scale factor $\theta_{2}(<0)$ to fit the drone-view with short distance. Note that as a global feature, the size of the last partition is constant.
@@ II Related Work
We briefly review related work, including cross-view geo-localization and part-based representation learning.
@@ II-A Cross-view Geo-localization
Benefiting from the powerful representation ability of neural networks, prevalent methods employ pre-trained backbone networks to extract deep features, thereby enabling cross-view image matching within a high-dimensional space. Numerous works [34, 35, 36, 37, 38, 39, 40, 41] have driven the development of the matching of ground and satellite views in terms of data and algorithms. For instance, CVUSA [36], CVACT [37] and Vigor [40] are extensively utilized as benchmark datasets to evaluate algorithm performance. Deuser et al. [38] designed GPS-sampling and dynamic similarity sampling to select hard negatives, leading to superior results on several datasets. To release the cost of collecting precise pairwise data, Li et al. [42] proposed a unsupervised framework to utilize unlabeled data in ground-view geo-localization. In addition, image synthesis methods [43, 44, 45] minimizing appearance differences of cross-views, can significantly improve the matching accuracy.
@@
Figure 3: The comparison of our method with typical partition strategies, including soft-partition strategy (a) FSRA [19] and hard-partition strategy (b) LPN [24]. The first and third rows are drone-view images with long and short distance, while the second row is the matched satellite image with a single scale. For different partition strategies, matched part-level image representations are placed in the same column.
@@
With the growing popularity of UAV devices, several datasets [46, 47, 48] introduced drone-view images to support for the UVGL task. Zheng et al. [10] collected the University-1652 dataset, which contains pairs of ground-views, drone-views and satellite-views. Notably, Zheng [10] introduced variations in both the shooting angle and shooting distance of drone-views, thus enhancing the diversity of University-1652. To investigate effects of drone-views at various heights on geo-localization, Zhu et al. [46] developed SUES-200 comprising UAV images acquired at four altitudes. DenseUAV [47], as the pioneering low-altitude urban scene dataset, was tailored for the UAV self-positioning task (i.e, geo-localization within continuous local areas). Focusing on inaccurate alignment between query images and corresponding labels, GTA-UVA [48] dataset introduced semi-positive samples to simulate imperfect matching. Despite viewpoint changes compared to ground-view geo-localization, mainstream solutions [10, 11, 12, 13, 49, 50, 51, 52, 53, 54] converged upon a similar principle: mapping multi-view images into a shared high-dimensional space, subsequently facilitating image matching through similarity calculation. Several studies [19, 26, 55] proved that stronger backbone networks can significantly improve the accuracy of UVGL. Sun et al. [52] designed a specific feature extractor named F3-Net to capture both local and global information. RK-Net [12] inserted unit subtraction attention modules inside the backbone to detect representative keypoints, yielding robust features against viewpoints. Considering the color discrepancy between drone-views and satellite-views, recent methods [56, 57, 58] adjusted the color distribution of drone-views to reduce domain gaps. Some efforts [20, 51, 59] improved the optimization function in the training phase to facilitate domain alignment of cross-view images. To obtain robust semantic representations, plentiful partition learning strategies are also introduced, which will be detailed in next subsection. In this paper, we propose a UAV-view geo-localization algorithm applicable to scenes with inconsistent cross-view scales. To meet the requirements of scale inconsistency and scene diversity, we conduct experiments on University-1652 [10].
@@ II-B Part-based Representation Learning
Fine-grained features guide models to capture more comprehensive information, which has been proven effective in many fields, such as image restoration [60, 61, 62], recognition [63, 64, 65] and segmentation [66, 67, 68]. Inspired by classical part-based descriptors [69, 70, 71], numerous works in person re-identification [72, 73, 74, 75, 76, 77, 78] developed local feature representation by introducing human body structure information. For instance, Spindle-Net [72] leveraged body joints to capture consistent body features from different images, thus enabling the alignment of macro- and micro-body features across images. To eliminate the dependence on skeleton points, researchers further proposed coarse partition strategies rooted in the vertical distribution of the human body [76, 77, 78].
@@
Previous research has laid the foundation for partition learning in UVGL, which can be divided into two categories based on the pattern of feature segmentation: soft-partition learning [19, 20, 21, 22] and hard-partition learning [24, 25, 26, 27, 28, 29]. Soft-partition strategies, represented by FSRA [19], leverage feature values to distinguish categories of semantic information. Subsequently, TransFG [20] adopts the second-order gradient of features as the partition index, yielding a decent performance boost. FSRA and TransFG divided features equally based on the setting that each part has equal area, resulting in misalignment of cross-view partition pairs in both large-scale and small-scale scenes. To address this issue, GeoFormer [21] utilized a k-modes clustering algorithm to divide features adaptively. However, the above soft-partition strategies rely on the transformer framework [79]. Drawing upon the similarity between drone- and satellite-views, some methods employ fixed templates to generate fine-grained features, which are defined as hard-partition strategies. LPN [24] designed non-overlapping square-ring templates to segment features, thus mining contextual information. Based on LPN, several studies [31, 32, 33, 29, 28] optimized the shape and number of templates to further facilitate discriminative representation. Aiming at the problem of position offset, SDPL [55] proposed a shifting-fusion strategy, which adjusts the segmentation center to align partitions with non-centered targets. The hard partition strategies exhibit commendable accuracy and compatibility, while being sensitive to cross-view scale variation. Recently, Safe-Net [30] proposed a saliency-guided self-adaptive partition strategy, which achieves preferable results than previous methods in scale inconsistency scenarios. However, Safe-Net [30] also produces significant performance degradation in extreme scenarios. We intend to design an robust algorithm against scale variations for UVGL.
@@
To visually reveal the dependence of partition learning on cross-view scales, we illustrate two representative methods (FSRA [19] and LPN [24]) for dealing with long distance scenes and short distance scenes, as shown in Fig. 3. Limited to the setting where the partition area is the same, FSRA divided features of different categories into the same partition, resulting in feature misalignment. LPN employed fixed templates to generate fine-grained features. Notably, in scenarios where cross-view scales are inconsistent, partition pairs contain misaligned semantic content. In this paper, we study a part-based representation learning for UVGL. In comparison with soft-partition strategies, our method achieves superior performance and compatibility, whereas, in contrast to hard-partition strategies, it exhibits enhanced robustness against scale variations.
III Methods
As depicted in Fig. 4, our DGDPL framework consists of three stages: feature extraction, distance guided dynamic partition learning and classification supervision. In the feature extraction stage, a backbone network maps drone- and satellite-view images to high-level feature maps. Then, we introduce distance guided dynamic partition learning to generate fine-grained features with consistent content. We also propose a saliency-guided refinement strategy to identify differences within part-level features. Finally, we adopt a classifier to predict the category of features obtained from different views, followed by using cross-entropy loss to optimize the model.

### v3

# Scale-adaptive UAV Geo-Localization via Height-aware Partition Learning

@@ Abstract
UAV Geo-Localization faces significant challenges due to the drastic appearance discrepancy between drone-captured images and satellite views. Existing methods typically assume a consistent scaling factor across views and rely on predefined partition alignment to extract viewpoint-invariant representations through part-level feature construction. However, this scaling assumption often fails in real-world scenarios, where variations in drone flight states lead to scale mismatches between cross-view images, resulting in severe performance degradation. To address this issue, we propose a scale-adaptive partition learning framework that leverages known drone flight height to predict scale factors and dynamically adjust feature extraction. Our key contribution is a height-aware adjustment strategy, which calculates the relative height ratio between drone and satellite views, dynamically adjusting partition sizes to explicitly align semantic information between partition pairs. This strategy is integrated into a Scale-adaptive Local Partition Network (SaLPN), building upon an existing square partition strategy to extract both fine-grained and global features. Additionally, we propose a saliency-guided refinement strategy to enhance part-level features, further improving retrieval accuracy. Extensive experiments validate that our height-aware, scale-adaptive approach achieves state-of-the-art geo-localization accuracy in various scale-inconsistent scenarios and exhibits strong robustness against scale variations. The code will be made publicly available.
@@
image retrieval, geo-localization, dynamic partition, relative height ratio, scale variations
@@ I Introduction
Unmanned Aerial Vehicles (UAVs), also known as drones, have gained popularity due to their ability to efficiently capture data with few occlusions and rich content. Differ from ground-views, UAVs can easily acquire multi-scale images by changing the flight altitude and camera parameters. Benefiting from these advantages, UAVs play an irreplaceable role in various fields, including agricultural operations [1], automatic driving [2] and aerial photography [3]. Regardless of the applications, precise positioning and navigation of drones are indispensable, which mainly rely on a global navigation satellite system (GNSS) and high-quality communication environment [4]. In scenarios with weak or failed signals, GNSS will incorrectly determine the geographic localization of drones. Therefore, vision-based UAV-view geo-localization (UVGL) develops a meaningful hot research, aiming to release the dependence on GNSS. As the key of this task, cross-view matching algorithms calculates the similarity between drone- and satellite-view images to achieve image matching. However, due to the appearance difference between drone-views and satellite-views, traditional image matching algorithms [5, 6, 7] are imprecise.
@@
Figure 1: The simplified diagram of our research motivation. Given drone-view images captured from different heights (high-yellow box, middle-green box, low-blue box) and the corresponding satellite-view image (right), we propose a scale-adaptive partition learning strategy. This strategy dynamically adjusts to different altitudes by eliminating redundant background and expanding the field of view (FoV) to enhance spatial alignment. As shown, the extended/indented partition pairs exhibit more semantically consistent content, facilitating subsequential representation learning.
@@
With the rapid development of deep learning, current methods [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] utilize pre-trained backbone networks to extract high-level semantic representations and perform image matching based on feature distances, achieving superior accuracy than non-learning-based algorithms. To obtain robust discriminative representations, plentiful partition learning strategies have been proposed, which construct fine-grained features by segmenting high-level features to fully exploit contextual information. Depending on the pattern of feature segmentation, partition learning can be divided into two categories: soft-partition learning [20, 21, 22, 23, 24] and hard-partition learning [25, 26, 27, 28, 29, 30]. The former classifies features based on feature values to distinguish the categories of discriminative information, so the shape of partitions is uncontrollable. FSRA [20], as a representative of soft-partition learning, confirms that fine-grained features have a consistent distribution with image content and can boost model performance. In contrast, inspired by the distribution similarity between drone- and satellite-views, hard-partition strategies set predefined templates with fixed shape to divide high-level features. As the pioneering hard-partition algorithm, LPN [25] designs non-overlapping square-ring templates to build fine-grained features. Subsequently, several studies [31, 32] further optimize the shape and number of templates to facilitate discriminative representation. Benefiting from the explicit utilization of contextual information that prevents the network from falling into a local optimum, hard-partition strategies reveal superior accuracy and compatibility than soft-partition strategies under same settings. Nevertheless, hard-partition strategies are more sensitive to spatial changes of input images. During the data collection process, the variations of drone flight state will lead to the scale mismatch of cross-view images. The predefined segmentation templates constrain misaligned semantic information to same parts, thus interfering with feature matching.
Figure 2: The proposed height-aware adjustment strategy ($N=3$ for illustration). Case (II): Assuming the drone and satellite has similar viewpoint height, we show the typical square partition process, i.e., the uniform partition. However, the assumption is not always hold, and thus, we consider a general form of partition. Case (I): When the drone is with higher flight height with extra backgrounds, the partition areas should be decreased based on the calculated scale factor $\theta_{1}$. Case (III): When the drone is with lower flight height with limited FoVs, the partition areas should be increased based on the calculated scale factor $\theta_{2}$.
@@
As shown in Fig. 1, we depict the limitation of hard-partition strategies in dealing with scale-misaligned image pairs. Taking LPN [25] as an example, the partition pairs contain similar semantic information in the case of similar cross-view scales (see drone-view with green box). When drone height changes, partition pairs of fixed shape contain misaligned semantic information, defined as imperfect matching (see drone-view with yellow or blue box). Following the perspective projection principle (“closer objects appear larger” in imaging), we argue that adjusting the partition size based on the shooting height of the drone can ensure the content consistency of partition pairs, thus boosting the matching of cross-view partition pairs. Concretely, for high-height drone-view, indent templates can reduce redundancy in the part-level representations, while expanded templates contributes to the content supplement of partitions for low-height drone-views. Imitating the process mentioned above, we propose a scale-adaptive partition learning framework named SaLPN, to enhance the content consistency of partition pairs when the shooting height of drone changes. Specifically, we design a square partition strategy (SPS) as shown in Case (II) of Fig. 2, which has two typical advantages: 1) multiple partitions constructed by SPS emphasize both fine-grained and global features, thereby exhibiting robustness against scale variations; 2) solid partitions have strong compatibility with subsequent algorithms. To promote the content consistency of partition pairs, we further propose a height-aware adjustment strategy (HAAS). We construct a scale factor $\theta$ based on the relative height ratios between drone and satellite views, and inject it into SPS to dynamically adjust the partition size of drone-view features. As shown in Case (I) and Case (III) of Fig. 2, HAAS controls variations of partitions, both in direction and degree. Furthermore, we design a saliency-guided refinement strategy (SGRS) to identify targets and environments based on feature activation degrees. SGRS maintains robustness to scale changes while improving model matching accuracy. Extensive experiments on University-1652 [8] and SUES-200 [33] show that SaLPN achieves state-of-the-art geo-localization accuracy in various scale-inconsistent scenarios and exhibits excellent robustness against scale variations. For instance, the average R@1 of SaLPN on multiple scale scenarios is 30.81% and 16.35% higher than that of Baseline with the instance loss [8] and LPN [25], respectively. Our main contributions are:
  • We introduce a new scale-adaptive partition learning framework that harness the pre-acquired drone height to explicitly addresses scale mismatches in cross-view scenarios. In particular, we dynamically adjust partition sizes based on relative height ratios between drone and satellite views, eliminating the conventional assumption of consistent scaling and improving semantic alignment.
  • We further develop a saliency-guided refinement strategy to enhance the discriminative ability of part-level features by merging the coordinate attention. Specifically, this strategy selectively fuses representations according to the heatmap and extracts the three-level features, i.e., global feature, salient feature and background feature.
@@
  • Extensive experiments show that, in comparison to both hard-partition and soft-partition strategies, our method achieves splendid retrieval performance, especially in scenes with inconsistent cross-view scales, which is common in the real-world drone flight.
@@ II Related Work
We briefly review related works, including cross-view geo-localization and part-based representation learning.
@@ II-A Cross-view Geo-localization
Benefiting from the powerful representation ability of neural networks, prevalent methods employ pre-trained backbone networks to extract deep features, thereby enabling cross-view image matching within a high-dimensional space. Numerous works [34, 35, 36, 37, 38, 39, 40] have driven the development of the matching of ground and satellite views in terms of data and algorithms. For instance, CVUSA [35], CVACT [36] and Vigor [39] are extensively utilized as benchmark datasets to evaluate algorithm performance. Deuser et al. [37] design GPS-sampling and dynamic similarity sampling to select hard negatives, leading to superior results on several datasets. To release the cost of collecting precise pairwise data, Li et al. [41] propose a unsupervised framework to utilize unlabeled data in ground-view geo-localization. In addition, image synthesis methods [42, 43, 44] minimizing appearance differences of cross-view images, can improve the matching accuracy.
@@
Figure 3: The comparison of our method with typical partition strategies, including (a) soft-partition strategy FSRA [20], (b) hard-partition strategy LPN [25] and (c) our height-aware method. The first and third rows are drone-view images with higher and lower heights, while the second row is the matched satellite image with a single scale. For different partition strategies, the corresponding part-level image representations are placed in the same column. We could observe that the proposed method has yielded a more consistent inter-part alignment at the same part level (i.e., every column).
@@
With the growing popularity of UAV devices, several datasets [33, 45, 46] introduce drone-view images to support for the UVGL task. Zheng et al. [8] collect the University-1652 dataset, which contains pairs of ground-views, drone-views and satellite-views. Notably, Zheng et al. [8] introduce variations in both the shooting angle and shooting height of drone-views, thus enhancing the diversity of University-1652. To investigate effects of drone-views at various heights on geo-localization, Zhu et al. [33] develop SUES-200 comprising drone images acquired at four altitudes. DenseUAV [45], as the pioneering low-altitude urban scene dataset, is tailored for the drone self-positioning task. Focusing on inaccurate alignment between query images and corresponding labels, GTA-UVA [46] introduces semi-positive samples to simulate imperfect matching. Despite viewpoint changes compared to ground-view geo-localization, mainstream solutions [8, 9, 10, 11, 47, 48, 49, 50, 51, 52] converge upon a similar principle: mapping multi-view images into a shared high-dimensional space, subsequently facilitating image matching through similarity calculation. Several studies [20, 26, 53] prove that stronger backbone networks can significantly improve the accuracy of UVGL. Sun et al. [50] design a specific feature extractor named F3-Net to capture both local and global information. RK-Net [10] inserts unit subtraction attention modules inside the backbone to detect representative keypoints, yielding robust features against viewpoints. Considering the color discrepancy between drone-views and satellite-views, recent methods [54, 55] adjust the color distribution of drone-views to reduce domain gaps. Some efforts [21, 49] improve the optimization function in the training phase to facilitate domain alignment of cross-view images. To obtain robust semantic representations, plentiful partition learning strategies are also introduced, which will be detailed in next subsection. In this paper, we propose a UAV-view geo-localization algorithm applicable to scenes with inconsistent cross-view scales. To meet the requirements of scale inconsistency and scene diversity, we conduct experiments on University-1652 [8].
@@ II-B Part-based Representation Learning
Fine-grained features guide models to capture more comprehensive information, which has been proven effective in many fields, such as image recognition [56, 57, 58] and segmentation [59, 60, 61]. Inspired by classical part-based descriptors [62, 63, 64], numerous works in person re-identification [65, 66, 67, 68, 69, 70] develop local feature representation by introducing human body structure information. For instance, OG-Net [70] leverages 3D body keypoints to capture consistent body features from 2D images, thus enabling the alignment of macro- and micro-body features across images. To eliminate the dependence on skeleton points, researchers further propose coarse partition strategies rooted in the vertical distribution of the human body [68, 69, 71].
@@
Previous research has laid the foundation for partition learning in UVGL, which can be divided into two categories based on the pattern of feature segmentation: soft-partition learning [20, 21, 22, 23] and hard-partition learning [25, 26, 27, 28, 29]. Soft-partition strategies, represented by FSRA [20], leverage feature values to distinguish categories of semantic information. Subsequently, TransFG [21] adopts the second-order gradient of features as the partition index, yielding a decent performance boost. FSRA and TransFG divide features equally based on the setting that each part has equal area, resulting in misalignment of cross-view partition pairs in both large-scale and small-scale scenes. To address this issue, GeoFormer [22] utilizes a k-modes clustering algorithm to divide features adaptively. However, the above soft-partition strategies rely on the transformer framework [72]. Drawing upon the similarity between drone- and satellite-views, some methods employ fixed templates to generate fine-grained features, which are defined as hard-partition strategies. LPN [25] designs non-overlapping square-ring templates to segment features, thus mining contextual information. Based on LPN, several studies [31, 32, 29, 28] optimize the shape and number of templates to further facilitate discriminative representation. Aiming at the problem of position offset, SDPL [53] proposes a shifting-fusion strategy, which adjusts the segmentation center to align partitions with non-centered targets. The hard partition strategies exhibit commendable accuracy and compatibility, while being sensitive to cross-view scale variation. Recently, Safe-Net [30] proposes a saliency-guided self-adaptive partition strategy, which achieves preferable results than previous methods in scale inconsistency scenarios. However, Safe-Net [30] also produces significant performance degradation in extreme scenarios. We intend to design an robust algorithm against scale variations for UVGL.
@@
To visually reveal the dependence of partition learning on cross-view scales, we illustrate two representative methods (FSRA [20] and LPN [25]) for dealing with high-altitude scenes and low-altitude scenes, as shown in Fig. 3. Limited by the setting of equal partition area, FSRA divides features of different categories into the same partition, resulting in feature misalignment. LPN employs fixed templates to generate fine-grained features. Notably, in scenarios where cross-view scales are inconsistent, partition pairs contain misaligned semantic content. In this paper, we study a part-based representation learning for UVGL. In comparison with soft-partition strategies, our method achieves superior performance and compatibility, whereas, in contrast to hard-partition strategies, it exhibits enhanced robustness against scale variations.
III Methodology